OpenStudy (anonymous):

Which logarithmic equation has no solutions? a. ln(x)=1 b. ln(x+2)=ln(7-4x) c. ln(x-2)=1 d. ln(x-2)=ln(4x+7)

OpenStudy (anonymous):

D, solve for x and try to put in log

OpenStudy (anonymous):

log (-ive no.)= not defined

OpenStudy (kainui):

To undo a natural log (ln) just raise both sides to a power of e. ln(x)=1 e^(lnx)=e^1 x=e